• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Boeglin

    Log entry time 16:02:41 on June 5,2010

    Entry number 324727

    keyword=Shift summary

    8:00 production running current run 4299:
    8:37 end 4299: 78.5K muiltplets, 145 mC
    8:37 start run 4300:
    9:15 end of run: 53.8K
    Have compton dipole turned on
    went to empty target


    10:00 MCC still tuning for compton,
    turned heater on and cooling off

    10:30 beam back, heater off, cooler on
    10:40 ask for harp scan
    10:50 harp scans completed, goto Pb target
    11:00 take low current data to check traget
    11:20 ramp up to 50uA

    Start new production run: 4305

    1:30 Beam disappeared due to 2k cold box trip, expect longer down time
    stopped run 4307